OpenStudy (anonymous):

The Smiths plan to recarpet their family room, which measures 15 ft. by 20 ft. How many square yards of carpet are needed? does anyone know this

OpenStudy (anonymous):

33.33

myininaya (myininaya):

15ft*20ft=300ft^2 1 foot = 1/3 yard (1 foot)^2=(1/3 yard)^2 1 ft^2=1/9 yard^2 so we got 300 ft^2 but we want this in square yards so multiply both sides of 1 ft^2=1/9 yard^2 by 300 300 ft^2=300/9 yard^2 so the answer is 100/3 yard^2

OpenStudy (anonymous):

100/3yrds?